OpenStudy (anonymous):

Find the indefinite integral \[\int\limits_{}\tan^{-1}xdx = \int\limits_{}arctanxdx\]

OpenStudy (anonymous):

There we go, I got something ^.^

OpenStudy (anonymous):

writing \[u=\tan^{-1}x \] and \[v = x\] and using the integration by parts we have find, \[du = (1) / ( x^2 + 1 )\] and v = r

OpenStudy (anonymous):

so lets roll ^.^

OpenStudy (anonymous):

\[\int\limits_{}\tan^{-1}xdx =(\tan^{-1}x) - \int\limits_{}(r)( 1 / (1 + r^2))dr\] \[=x \tan^{-1}x - (1/2)\int\limits_{}(2xdx)/(1 + r^2)\] \[=x \tan^{-1}x - (1/2)\int\limits_{}dy/y\] setting \[y = r^2 + 1\] ... = \[x \tan^{-1}x - (1/2)\ln|y| + c \] =\[x \tan^{-1}x - (1/2)\ln(r^2 + 1) + c \] since \[y = r^2 + 1 > 0\]